i dont understand how this answer would come out as (-69)...... The problem is this , -33+(-7)=29+m . if you could please hellp me ASAP that would be great... thank you !
Answer by Alan3354(69443)

-33+(-7)=29+m
Add the -33 and the -7
-40 = 29+m
Subtract 29
-69 = m
m = -69
What else could it be?